Area bounded by the curve y=sin x between x=0 & x=360 is

Grade:Upto college level

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Area bounded by sinx in (0,360) is given by
\int_{0}^{2\pi }sinx=
2\int_{0}^{\pi }sinx=2(-cosx) in the limits (0,\pi )=-2(-2)=4
Hence the area is 4 sq.units
Thanks and regards
Shaik Aasif
askIITians faculty
SHAIK AASIF AHAMED
9 years ago
Hello student,
Area bounded by sinx in (0,360) is given by
\int_{0}^{2\pi }sinx=
2\int_{0}^{\pi }sinx=2(-cosx) in the limits (0,\pi )=-2(-2)=4
Hence the area is 4 sq.units
Thanks and regards
Shaik Aasif
askIITians faculty